# FlightPatchNet: Multi-Scale Patch Network with Differential Coding for Flight Trajectory Prediction

### v1

@@ 3.2 Multi-Scale Patch Network
Considering different temporal patterns prefer diverse time scales, the multi-scale patch network first utilizes a stack of $K$ patch mixer blocks to capture underlying temporal patterns from large scale to small scale. A large time scale can reflect the slow-varying flight trends, while a smaller scale can retain fine-grained local details. To further promote the collaboration of diverse temporal features, a multi-scale aggregator is introduced to consider the contributed scales and dominant variables. Such a multi-scale network equips our model with the powerful and complete temporal modeling capability, and helps preserve all kinds of multi-scale characteristics.
@@ 3.2.1 Patch Mixer Block
Only considering one single time step is insufficient for FTP task, since it contains limited semantic information and cannot accurately reflect the flight trajectory variations. Inspired by 2021An ; Yuqietal-2023-PatchTST , the trajectory representation $\mathbf{Z}\in\mathbb{R}^{C\times L}$ is segmented into several non-overlapping patches along the temporal dimension, generating a sequence of patches $\mathbf{Z}_{p}\in\mathbb{R}^{C\times P\times N}$, where $P$ is the length of each patch, $N$ represents the number of patches, and $N=\left\lceil\frac{L}{P}\right\rceil$. The patching process is formulated as:
@@
Based on the divided patches $\mathbf{Z}_{p}$, we utilize a patch encoder and decoder to capture temporal features in flight trajectories. Specifically, the patch encoder aims to capture the inter-patch features (i.e., the global correlations across patches) and intra-patch features (i.e., the local details within patches). After that, these features are reconstructed to the original dimension by the patch decoder. Due to the superiority of linear models for MTS chen2023tsmixer ; zeng2023transformers , the patch encoders and decoders are based on pure multi-layer perceptron (MLP) for temporal modeling.
@@
[…]
After such patch encoding process, the correlations between nearby time steps within patches and distant time steps across patches are finely explored. Then, we utilize a patch decoder to reconstruct the original sequence. A patch decoder comprises the same components as the encoder in a reverse order, which is formulated as follows:
@@ 3.2.2 Multi-Scale Aggregator
To enable the ability of more complete multi-scale modeling, we introduce a multi-scale aggregator to integrate different temporal patterns. It contains two components: scale fusion and channel fusion. Scale fusion can figure out critical time scales and capture the scale-wise correlations, while channel fusion can discover dominant variables effecting temporal variations and explore the inter-variable relationships. These two components work together to help the model learn a robust multi-scale representation and improve generalization ability.
Given the $K$ scale-specific temporal representations $\{\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K}\}$, we first stack them and rearrange the data to combine the three dimensions of channel size $C$, patch size $P$ and patch quantity $N$, resulting in $\mathbf{S}^{0}\in\mathbb{R}^{K\times(C\times L)}$, where $L=P\times N$. Then we apply MSA on the scale dimension $K$ to learn the importance of contributed time scales. This process is formulated as:
@@
$\mathbf{S}^{i}$ $=LayerNorm(\mathbf{S}^{i-1}+MSA(\mathbf{S}^{i-1}),i=1,\dots,l$
@@
where $\mathbf{S}^{l}$ is the final multi-scale fusion representation within variables. Inspired by liu2023itransformer , we consider each variable as a token and apply MSA to explore dependencies between different variables. We first reshape the $\mathbf{S}^{l}$ to get $\mathbf{C}^{0}$ $\in\mathbb{R}^{C\times(K\times L)}$ and perform multi-head self-attention on the channel dimension $C$ to identify dominant variables. This process is simply formulated as follows:
@@
[…]
@@ 3.3 Direct Multi-Step Prediction
[…]
@@
Finally, all the predictor results are aggregated to generate the final prediction, which can enhance the stability and generalization of our model.
@@ 4.1 Dataset and Experimental Setup
To evaluate the performance of FlightPatchNet, we conduct extensive experiments on ADS-B data provided by OpenSky  [^1] https://opensky-network.org/datasets/states/ from 2020 to 2022. In this paper, six key attributes are extracted from the original data, including longitude, latitude, altitude, and velocity in x, y, z dimensions. The dataset is divided into three parts for training, validation, and testing with a ratio of 8:1:1.
@@
We compare our model with five competitive models, including four IMS-based models: LSTM LSTM8489734 , Bi-LSTM Sahadevan , CNN-LSTM CNN-LSTM9145522 , FlightBERT Guo2023FlightBERT ; one DMS-based model: FlightBERT++ Guo2023FlightBERT++ . These models have covered mainstream deep learning architectures, including Transformer(FlightBERT, FlightBERT++), CNN(CNN-LSTM), and RNN(LSTM, Bi-LSTM, CNN-LSTM), which help to provide a comprehensive comparison. For fairness, all the models are following the same experimental setup with lookback window $L=60$ and prediction horizon $T\in\{1,3,9,15\}$. Our model is trained with MSE loss, using the Adam optimizer kingma2014adam . We adopt the Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as evaluation metrics. More details about dataset, baselines, implementation and hyper-parameters are shown in Appendix A.
@@ 4.2 Main results
Comprehensive flight prediction results are demonstrated in Table 1 (see Appendix B.2 for error bar). FlightPatchNet achieves the most outstanding performance across various prediction lengths for longitude and latitude in terms of both MAE and RMSE, while does not achieve the optimal results for altitude compared with other strong baselines such as FlightBERT++. For simplification, we consider prediction horizon $T=15$ and compare our model with the second best. FlightPatchNet achieves an overall 18.62% reduction on MAE and 49.63% reduction on RMSE for longitude, and 35.31% reduction on MAE and 44.80% reduction on RMSE for latitude. For the prediction performance of altitude, FlightBERT++ outperforms our model by 45.51 meters reduction on MAE but has a large RMSE which may caused by high-bit errors in the prediction. FlightPatchNet obtains the smallest RMSE results for all variables, indicating that our model can provide a more robust and stable prediction. Furthermore, as the prediction horizon increases, DMS-based models greatly outperform IMS-based models which suffer from serious performance degradation due to error accumulation. Note that longitude, latitude and altitude are combined together to determine the position of an aircraft. Poor prediction performance on any variable is intolerable and meaningless for short-term FTP task. Thus, FlightPatchNet achieves the most competitive performance in general.
@@
Model Metric Lon() Lat() Alt(m)
@@
IMS LSTM MAE 0.00056 0.00427 0.01747 0.03132 0.00049 0.00493 0.02116 0.03717 92.27 159.30 549.55 882.86
RMSE 0.00095 0.00691 0.02597 0.04578 0.00089 0.00740 0.02956 0.05143 142.05 233.39 763.84 768.45
Bi-LSTM MAE 0.00155 0.00747 0.02319 0.03890 0.00137 0.00824 0.02711 0.04404 432.50 761.50 1648.68 2006.21
RMSE 0.00202 0.01124 0.03387 0.05532 0.00181 0.01142 0.03639 0.05982 563.74 953.37 2132.91 2420.74
CNN-LSTM MAE 0.00139 0.00700 0.02282 0.04149 0.00131 0.00801 0.02623 0.05139 520.03 746.67 1569.68 1136.80
RMSE 0.00240 0.01033 0.03263 0.05981 0.00212 0.01130 0.03559 0.07353 1176.96 926.40 1936.63 1658.53
FlightBERT MAE 0.00123 0.00241 0.01162 0.02407 0.00088 0.00158 0.00963 0.01238 24.67 35.67 78.58 134.29
RMSE 0.00241 0.00526 0.02189 0.03969 0.00154 0.00286 0.01904 0.03093 234.17 272.59 384.22 462.28
DMS FlightBERT++ MAE 0.00173 0.00317 0.00871 0.01187 0.00085 0.00210 0.00612 0.01048 9.39 21.89 47.84 78.46
RMSE 0.00360 0.00659 0.01846 0.03131 0.00148 0.00425 0.00959 0.02127 175.29 167.16 327.93 384.18
FlightPatchNet (Ours) MAE 0.00048 0.00153 0.00546 0.00966 0.00032 0.00105 0.00381 0.00678 13.34 32.65 78.57 123.97
RMSE 0.00087 0.00233 0.00885 0.01577 0.00064 0.00175 0.00652 0.01174 123.78 121.48 174.63 244.34
4.3 Effectiveness of Differential Coding
The results in Table 2 show that using differential coding for longitude and latitude can significantly improve their prediction performance but slightly decrease the accuracy of altitude. The differential coding can reveal the temporal variations of longitude and latitude, which helps the temporal modeling in flight trajectories. However, the variations of altitude in original series may come from unexpected noise. FlightPatchNet has strong modeling capacity for temporal variations and tends to focus on the noise points during altitude prediction, leading to a large bias towards the ground truth. Further analysis are presented in Appendix C.
@@ 4.3 Effectiveness of Differential Coding
Models Diff Lon() Lat() Alt(m)
MAE RMSE MAE RMSE MAE RMSE
LSTM $\checkmark$ 0.03132 0.04578 0.03717 0.05143 882.86 1205.78
$\times$ 0.82230 1.20424 0.12008 2.44136 768.45 1053.21
Bi-LSTM $\checkmark$ 0.03890 0.05532 0.04404 0.05982 2006.21 2420.74
$\times$ 1.71433 2.43607 0.19014 0.27621 2091.19 2665.51
CNN-LSTM $\checkmark$ 0.04149 0.05981 0.05139 0.07353 1136.80 1658.53
$\times$ 8.59512 23.07600 1.95957 8.15418 1638.02 2113.49
FlightPatchNet (ours) $\checkmark$ 0.00966 0.01577 0.00678 0.01174 123.97 244.34
$\times$ 0.19348 0.26243 0.05385 0.07457 60.63 169.57
4.4 Effectiveness of Multi Scales
Table 3: The flight trajectory prediction results of single scales and multi scales. The best results are highlighted in bold.
Lon() Lat() Alt(m)
Patch Size Metric 1 3 9 15 1 3 9 15 1 3 9 15
2 MAE 0.00050 0.00156 0.00564 0.01038 0.00034 0.00105 0.00388 0.00711 16.50 23.39 93.28 106.63
RMSE 0.00091 0.00248 0.00907 0.01675 0.00065 0.00176 0.00654 0.01218 122.41 105.19 202.30 225.08
6 MAE 0.00052 0.00158 0.00583 0.01027 0.00036 0.00106 0.00402 0.00715 12.72 22.57 120.51 127.15
RMSE 0.00091 0.00251 0.00937 0.01656 0.00067 0.00178 0.00670 0.01227 106.48 106.58 226.89 247.28
10 MAE 0.00049 0.00162 0.00570 0.01019 0.00033 0.00108 0.00382 0.00700 12.16 35.77 107.07 178.91
RMSE 0.00088 0.00256 0.00912 0.01652 0.00064 0.00180 0.00654 0.01201 126.31 101.96 216.02 306.40
20 MAE 0.00054 0.00163 0.00562 0.01032 0.00036 0.00110 0.00391 0.00707 15.58 25.57 95.56 152.48
RMSE 0.00091 0.00257 0.00907 0.01660 0.00066 0.00182 0.00659 0.01211 99.66 106.19 202.99 286.20
30 MAE 0.00050 0.00161 0.00580 0.01065 0.00033 0.00108 0.00399 0.00716 10.35 24.75 75.23 144.55
RMSE 0.00090 0.00256 0.00935 0.01719 0.00065 0.00180 0.00676 0.01220 104.48 99.18 182.47 273.72
30,20,10,6,2 MAE 0.00048 0.00153 0.00546 0.00966 0.00032 0.00105 0.00381 0.00678 13.34 32.65 78.57 123.97
RMSE 0.00087 0.00233 0.00885 0.01577 0.00064 0.00175 0.00652 0.01174 129.65 121.78 174.63 244.34
To investigate the effect of multi-scale modeling, we conduct experiments on single scale for {2,6,10,20,30}. The results in Table 3 illustrate the critical contribution of multi scales to our model. We can observe that different variables prefer distinct time scales. For example, using patch size 10 can obtain the second best prediction performance on longitude and latitude but the worst performance on altitude when prediction horizon $T=15$. This indicates that longitude, latitude and altitude have distinct temporal patterns, and different scales can extract diverse complementary features, which can be effectively leveraged to obtain competitive and robust prediction results.

### v3

@@ 3.2 Multi-Scale Patch Network
Considering different temporal patterns prefer diverse time scales, the multi-scale patch network first utilizes a stack of $K$ patch mixer blocks to capture underlying temporal patterns from large scale to small scale. A large time scale can reflect the slow-varying flight trends, while a smaller scale can retain fine-grained local details. To further promote the collaboration of diverse temporal features, a multi-scale aggregator is introduced to consider the contributed scales and dominant variables. Such a multi-scale network equips our model with powerful and complete temporal modeling capability and helps preserve all kinds of multi-scale characteristics.
@@ 3.2.1 Patch Mixer Block
Only considering one single time step is insufficient for the FTP task, since it contains limited semantic information and cannot accurately reflect the flight trajectory variations. Inspired by PatchTST [Nie et al., 2023], the trajectory representation $\mathbf{Z}\in\mathbb{R}^{C\times L}$ is segmented into several non-overlapping patches along the temporal dimension, generating a sequence of patches $\mathbf{Z}_{p}\in\mathbb{R}^{C\times P\times N}$, where $P$ is the length of each patch, $N$ represents the number of patches, and $N=\left\lceil\frac{L}{P}\right\rceil$. The patching process is formulated as:
@@
Based on the divided patches $\mathbf{Z}_{p}$, we utilize a patch encoder and decoder to capture temporal features in flight trajectories. Specifically, the patch encoder aims to capture the inter-patch features (i.e., the global correlations across patches) and intra-patch features (i.e., the local details within patches). After that, these features are reconstructed to the original dimension by the patch decoder. Due to the superiority of linear models for MTS [Chen et al., 2023b, Zeng et al., 2023], the patch encoders and decoders are based on pure multi-layer perceptron (MLP) for temporal modeling.
@@
$\mathbf{E}=\mathit{Linear}(\mathbf{N}_{intra}^{\top})$ (6)
After such a patch encoding process, the correlations between nearby time steps within patches and distant time steps across patches are finely explored. Then, we utilize a patch decoder to reconstruct the original sequence. A patch decoder comprises the same components as the encoder in a reverse order, which is formulated as follows:
@@ 3.2.2 Multi-Scale Aggregator
To enable the ability of more complete multi-scale modeling, we introduce a multi-scale aggregator to integrate different temporal patterns. It contains two components: scale fusion and channel fusion. Scale fusion can figure out critical time scales and capture the scale-wise correlations, while channel fusion can discover dominant variables affecting temporal variations and explore the inter-variable relationships. These two components work together to help the model learn a robust multi-scale representation and improve generalization ability. Given the $K$ scale-specific temporal representations $\{\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K}\}$, we first stack them and rearrange the data to combine the three dimensions of channel size $C$, patch size $P$ and patch quantity $N$, resulting in $\mathbf{S}^{0}\in\mathbb{R}^{K\times(C\times L)}$, where $L=P\times N$. Then we apply MSA on the scale dimension $K$ to learn the importance of contributed time scales. This process is formulated as:
@@
$\mathbf{S}^{i}$ $=LayerNorm(\mathbf{S}^{i-1}+MSA(\mathbf{S}^{i-1})$
@@
where $\mathbf{S}^{l}$ is the final multi-scale fusion representation within variables. Inspired by iTransformer [Liu et al., 2023], we consider each variable as a token and apply MSA to explore dependencies between different variables. We first reshape the $\mathbf{S}^{l}$ to get $\mathbf{C}^{0}$ $\in\mathbb{R}^{C\times(K\times L)}$ and perform multi-head self-attention on the channel dimension $C$ to identify dominant variables. This process is simply formulated as follows:
@@
$\mathbf{C}^{i}$ $=LayerNorm(\mathbf{C}^{i-1}+\mathit{MSA}(\mathbf{C}^{i-1}))$
$\mathbf{C}^{i}$ $=\mathit{LayerNorm}(\mathbf{C}^{i}+FC(\mathbf{C}^{i}),i=1,\dots,l$
@@ 3.3 Direct Multi-Step Prediction
We ensemble $K$ predictors to directly obtain the future flight trajectory series, which can exploit complementary information from different temporal patterns. The objective of our model is to predict the differential values of longitude and latitude relative to the last observation, and the raw absolute values of altitude, i.e., $\hat{\mathbf{Y}}=\left\{\hat{\mathbf{x}}_{L+1},...,\hat{\mathbf{x}}_{L+T}\right\}$, where $\hat{\mathbf{x}}_{L+i}=(\hat{\Delta}^{\mathit{Lon}}(L+i,L),\hat{\Delta}^{\mathit{Lat}}(L+i,L),\hat{Alt}_{L+i})^{\top}$ for $i=1,\dots,T$. We split the final multi-scale representation $\mathbf{H}\in\mathbb{R}^{C\times L\times K}$ into a sequence $\left\{\mathbf{H}_{*,1},\mathbf{H}_{*,2},\dots,\mathbf{H}_{*,K}\right\}$, where $\mathbf{H}_{*,i}\in\mathbb{R}^{C\times L}$ for $i=1,\dots,K$, and feed each $\mathbf{H}_{*,i}$ to a predictor. Each predictor has two MLPs. The first $MLP_{C_{i}}$ transforms the input channel $C$ into the output channel $C^{\prime}$, and the second $MLP_{T_{i}}$ projects the historical input sequence $L$ to the prediction horizon $T$.
@@
Finally, all the predictor results are aggregated to reconstruct the final prediction trajectory according to Equation (1), which can enhance the stability and generalization of our model.
@@ 4.1 Dataset and Experimental Setup
To evaluate the performance of FlightPatchNet, we conduct extensive experiments on ADS-B data provided by OpenSky  [^1] https://opensky-network.org/datasets/states/ from 2020 to 2022. In this paper, six key attributes are extracted from the original data, including longitude, latitude, altitude, and velocity in x, y, z dimensions. The dataset is chronologically divided into three parts for training, validation, and testing with a ratio of 8:1:1.
@@
We compare our model with ten competitive models, including five IMS-based models: LSTM [Shi et al., 2018], CNN-LSTM [Ma and Tian, 2020], Bi-LSTM [Sahadevan et al., 2022], FlightBERT [Guo et al., 2023], WTFTP [Zhang et al., 2023]; five DMS-based model: FlightBERT++ [Guo et al., 2024], TimeMixer [Wang et al., 2024], TimesNet [Wu et al., 2023a], MICN [Wang et al., 2023], Pathformer [Chen et al., 2024]. These models have covered mainstream deep learning architectures, including Transformer (FlightBERT, FlightBERT++, Pathformer), CNN (CNN-LSTM, TimesNet, MICN), RNN (LSTM, Bi-LSTM, CNN-LSTM, WTFTP) and MLP (TimeMixer), which helps to provide a comprehensive comparison. For fairness, all the models follow the same experimental setup with lookback window $L=60$ and prediction horizon $T\in\{1,3,9,15\}$. Our model is trained with MSE loss, using the Adam optimizer. We adopt the Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as evaluation metrics. More details about the dataset, baselines, implementation and hyper-parameters are shown in Appendix A.
@@ 4.2 Main Results
Comprehensive flight prediction results are demonstrated in Table 1 (see Appendix B.2 for error bar). FlightPatchNet achieves the most outstanding performance across various prediction lengths for longitude and latitude in terms of both MAE and RMSE, while it does not achieve the optimal results for altitude compared with other strong baselines such as FlightBERT++. For simplification, we consider prediction horizon $T=15$ and compare our model with the second best. FlightPatchNet achieves an overall 18.62% reduction on MAE and 41.29% reduction on RMSE for longitude, and 35.31% reduction on MAE and 44.80% reduction on RMSE for latitude. For the prediction performance of altitude, FlightBERT++ outperforms our model by 45.51 meters reduction on MAE but has a large RMSE which may caused by high-bit errors in the prediction. FlightPatchNet obtains the smallest RMSE results for all variables, indicating that our model can provide a more robust and stable prediction. Compared with the most promising multi-scale modeling MTS prediction methods, including the pure MLP-based model TimeMixer, the CNN-based methods TimesNet and MICN and the Transformer-based method Pathformer, FlightPatchNet achieves superior prediction performance. These existing multivariate time-series forecasting methods typically decompose time series into seasonal and trend components, and primarily focus on periodic modeling. However, short-term flight trajectories do not exhibit obvious periodic patterns. The trend features in altitude and the temporal variations in longitude and latitude deserve more attention. Furthermore, as the prediction horizon increases, IMS-based models suffer from serious performance degradation due to error accumulation.
@@
Model Metric Lon($0.00001^{\circ}$) Lat($0.00001^{\circ}$) Alt(m)
@@
IMS LSTM MAE 56 427 1747 3132 49 493 2116 3717 92.27 159.30 549.55 882.86
RMSE 95 691 2597 4578 89 740 2956 5143 142.05 233.39 763.84 768.45
Bi-LSTM MAE 155 747 2319 3890 137 824 2711 4404 432.50 761.50 1648.68 2006.21
RMSE 202 1124 3387 5532 181 1142 3639 5982 563.74 953.37 2132.91 2420.74
CNN-LSTM MAE 139 700 2282 4149 131 801 2623 5139 520.03 746.67 1569.68 1136.80
RMSE 240 1033 3263 5981 212 1130 3559 7353 1176.96 926.40 1936.63 1658.53
WTFTP MAE 175 1484 2002 2657 112 1169 1586 2110 145.02 230.49 588.44 957.41
RMSE 218 1905 2606 3531 171 1739 2328 3124 415.13 497.24 1021.52 1583.38
FlightBERT MAE 123 241 1162 2407 88 158 963 1238 24.67 35.67 78.58 134.29
RMSE 241 526 2189 3969 154 286 1904 3093 234.17 272.59 384.22 462.28
DMS FlightBERT++ MAE 173 317 871 1187 85 210 612 1048 9.39 21.89 47.84 78.46
RMSE 360 659 1846 3131 148 425 959 2127 175.29 167.16 327.93 384.18
TimeMixer MAE 67 765 3100 5581 42 422 1679 3028 21.18 50.02 119.80 157.47
RMSE 115 1466 5517 9698 76 789 2976 5318 109.57 145.55 281.31 374.82
TimesNet MAE 73 1383 5459 9421 46 753 2981 5086 36.41 79.42 178.49 257.39
RMSE 124 2281 8377 14420 83 1237 4557 7796 154.37 184.36 392.23 543.37
MICN MAE 69 680 2235 3912 40 384 1296 2256 48.28 544.35 1992.86 3431.19
RMSE 116 1133 3831 6598 73 633 2167 3704 112.31 945.57 3270.77 5612.64
Pathformer MAE 52 232 1374 1914 45 232 863 1806 42.03 47.65 141.08 259.53
RMSE 89 373 2227 2686 83 373 1346 2570 114.37 136.59 406.37 645.02
FlightPatchNet
MAE 48 153 546 966 32 105 381 678 13.34 32.65 78.57 123.97
RMSE 87 233 885 1577 64 175 652 1174 123.78 121.48 174.63 244.34
Visualization of FlightPatchNet Predictions
Figure 4: Visualization of the ground truth and predictions of FlightPatchNet when the prediction horizon $T=15$ and look-back window size $L=60$. The series of altitude are in original data while those of longitude and latitude are in differential data, all denoted by meters.
Visualization of FlightPatchNet Predictions for Altitude
Figure 5: Visualization of ground truth for altitude and predictions of FlightPatchNet when the prediction horizon $T=15$ and look-back window size $L=60$, all denoted by meters.
We present the visualization of FlightPatchNet predictions and ground truth for the altitude in Figure 5. As shown in Figure 5, when the series of altitudes are relatively smooth and stationary with obvious global trends, FlightPatchNet can effectively capture these trends and make accurate predictions. When the series suffers from many change points caused by frequent abrupt fluctuations, as depicted in Figure 5 and  5, FlightPatchNet tends to focus more on the irregular change points during prediction, leading to a large deviation from the ground truth. As a result, FlightPatchNet struggles to capture the real temporal variations in altitude and fails to provide accurate predictions.
Effectiveness of Differential Coding
The results in Table 2 show that using differential coding for longitude and latitude can significantly improve their prediction performance but slightly decrease the accuracy of altitude. The differential coding can reveal the temporal variations of longitude and latitude, which helps the temporal modeling in flight trajectories. However, the variations of altitude in the original series may come from unexpected noise. FlightPatchNet has a strong modeling capacity for temporal variations and tends to focus more on the noise points during altitude prediction, leading to a large bias towards the ground truth.
@@
Models Diff Metric Lon() Lat() Alt(m)
LSTM $\checkmark$ MAE 0.03132 0.03717 883
RMSE 0.04578 0.05143 1206
$\times$ MAE 0.82230 0.12008 769
RMSE 1.20424 2.44136 1053
Bi-LSTM $\checkmark$ MAE 0.03890 0.04404 2006
RMSE 0.05532 0.05982 2421
$\times$ MAE 1.71433 0.19014 2091
RMSE 2.43607 0.27621 2666
CNN-LSTM $\checkmark$ MAE 0.04149 0.05139 1137
RMSE 0.05981 0.07353 1659
$\times$ MAE 8.59512 1.95957 1638
RMSE 23.07600 8.15418 2114
FlightPatchNet
$\checkmark$ MAE 0.00966 0.00678 124
RMSE 0.01577 0.01174 244
$\times$ MAE 0.19348 0.05385 61
RMSE 0.26243 0.07457 170
Effectiveness of Multi Scales
Figure 6: MAE and RMSE of FlightPatchNet for longitude with single scale and multi scales ($L=60$).
Figure 7: MAE and RMSE of FlightPatchNet for latitude with single scale and multi scales ($L=60$).
Figure 8: MAE and RMSE of FlightPatchNet for altitude with single scale and multi scales ($L=60$).
To investigate the effect of multi-scale modeling, we conduct experiments on single scale for {2,6,10,20,30}. The results in Figure 6, 7 and 8 illustrate the critical contribution of multi scales to our model. We observe that different variables prefer distinct time scales. For instance, a patch size of 10 obtains the second-best prediction performance on longitude and latitude but the worst performance on altitude when prediction horizon $T=15$. This indicates that longitude, latitude and altitude exhibit distinct temporal patterns, and different scales can extract diverse complementary features, which can be effectively leveraged to obtain competitive and robust prediction performance.